# FIRE [Uncaptioned image]: Flexible Integration of Data Quality Ratings for Effective Pretraining

### v2

@@ Abstract
Selecting high-quality data can significantly improve the pretraining efficiency of large language models (LLMs). Existing methods generally rely on heuristic techniques and single-quality signals, limiting their ability to evaluate data quality comprehensively. In this work, we propose FIRE, a flexible and scalable framework for integrating multiple data quality raters, which allows for a comprehensive assessment of data quality across various dimensions. FIRE aligns multiple quality signals into a unified space, and integrates diverse data quality raters to provide a comprehensive quality signal for each data point. Further, we introduce a progressive data selection scheme based on FIRE that iteratively refines the selection of high-quality data points. Experiments on the SlimPajama dataset reveal that FIRE outperforms other data selection methods and significantly enhances the pretrained model across a wide range of downstream tasks, with a 2.9% average performance improvement over Random and reducing the FLOPs necessary to achieve a certain performance level by more than half.
@@
Liangyu Xu4,  Xuemiao Zhang1,4,  Feiyu Duan2,4thanks: Equal contribution.,  Sirui Wang3,4thanks: Corresponding author.,  Jingang Wang4,  Xunliang Cai4 1 Peking University 2 Beihang University 3 Tsinghua University 4 Meituan zhangxuemiao@pku.edu.cnduanfeiyu@buaa.edu.cn {xuliangyu02, wangsirui, wangjingang02, caixunliang}@meituan.com
@@ 1 Introduction
Figure 1: Few-shot downstream performance with respect to pretraining FLOPs for random, FIRE, and FIRE progressive.
@@
Extensive experiments on the SlimPajama Soboleva et al. (2023) dataset demonstrate that by applying integrated ratings from multiple raters, our method achieves superior results across a variety of downstream tasks. Figure 1 illustrates that FIRE significantly enhances the pretrained model, providing an average performance increase of 2.9% while also reducing the FLOPs needed to achieve a specific performance level by more than half. We summarize our main contributions as follows:
@@
(3) Extensive experiments on the SlimPajama dataset demonstrate that FIRE enhances the pretrained model’s performance by an average of 2.9% while reducing the FLOPs required to achieve a certain performance level by more than half.
@@ 2.1 Overview of the Method
We propose FIRE, a method that flexibly integrates multiple raters to comprehensively evaluate data quality. It involves two key processes: (a) Rating Alignment and (b) Rater Integration. Figure 2 illustrates the overall framework of FIRE.
@@
$I(i)=\mathbf{A}(i)^{T}(\mathbf{o}\odot\bm{\gamma})$ (1)
where $\mathbf{A}(i)$ is the vector of aligned ratings for data point $i$, $\mathbf{o}$ is the overall orthogonality vector, $\bm{\gamma}$ represents the intrinsic reliability vector of the raters, and $\odot$ denotes the element-wise product.
@@ 2.2 Rating Alignment
Uniformly sample a subset $\mathcal{D}_{r}$ from the pretraining data $\mathcal{D}_{t}$.
@@
where $\mathcal{D}_{ij}$ signifies the subset sampled from the $j$-th rating interval for rater $R_{i}$. The win rate $w_{ij}$ is the proportion of samples $x$ in $\mathcal{D}_{ij}$ that GPT-4o determines have higher quality than the comparison sample $y$ from $\mathcal{D}_{r}$. The prompt for GPT-4o is detailed in Appendix C.2.
@@ 2.3 Rater Integration
Suppose for raters $R_{1},R_{2},\ldots,R_{n}$, each rater corresponds to a standard basis vector $\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}$ in the quality space. The integrated quality vector $\mathbf{q}(i)$ of data point $i$ can be expressed as:
$\mathbf{q}(i)=\sum_{j=1}^{n}\gamma_{j}A_{j}(i)\mathbf{v}_{j}$ (3)
where $\gamma_{j}$ denotes the intrinsic reliability of rater $j$, $A_{k}(i)$ denotes the rating for data point $i$ from rater $R_{k}$ after alignment. Ideally, if $\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}$ form an orthogonal basis, it is reasonable to measure the overall quality of the data using the L1 norm of $\mathbf{q}(i)$, as it represents the sum of the scores of data point $i$ across various orthogonal quality dimensions. However, $\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}$ are not necessarily completely independent. There may be raters $R_{i}$ and $R_{j}$ with a correlation coefficient $\rho>0$ and directly adding the corresponding aligned ratings would increase the weight of a particular quality dimension. To mitigate this issue, we define $O(i,j)$ to quantify the orthogonality of two raters $i$ and $j$ (the formalization of $O(i,j)$ can be found in Appendix C.5). For a rater $R_{i}$, we apply the sum of its orthogonality with all other raters to weight its rating. If a rater is highly correlated with others, we use a lower orthogonality to penalize. So the integrated rating for data point $i$ can expressed as
$I(i)=\sum_{j=1}^{n}\gamma_{j}o_{j}A_{j}(i)$ (4)
@@
where $d$ is the damping factor and $k$ denotes the $k$-th iteration. Since the introduction of the damping factor aims to address the rank sinks problem, which is not present in our graph, we find it reasonable to set $d=1$. Assuming the number of iterations is $\alpha$, and normalizing the final result, our final formula becomes:
@@
Final version of the integrated rating for data point $i$ can be expressed as:
$I(i)=\mathbf{A}(i)^{T}(\mathbf{o}\odot\bm{\gamma})$ (9)
where $\mathbf{A}(i)=[A_{1}(i),A_{2}(i),\ldots,A_{n}(i)]^{T}$ is the vector of aligned ratings for data point $i$ from all raters, $\mathbf{o}$ is the overall orthogonality vector, $\bm{\gamma}=[\gamma_{1},\gamma_{2},\ldots,\gamma_{n}]^{T}$ represents the intrinsic reliability vector of the raters, and $\odot$ denotes the element-wise (Hadamard) product.
@@
12:     Reduce the data to $\eta\%$ of its original size
@@
3 Progressive Data Selection Through FIRE
@@
Specifically, as shown in Algorithm 1, we first calculate the integrated ratings for the data points in $\mathcal{D}_{t}$ based on FIRE, then sort the data points, but we don’t select them right away based on these ratings. It is reduced to $\eta\%$ of its original size using a decay factor $\eta$. Then, the data is partitioned into $n$ segments based on integrated ratings’ quantiles. Orthogonality is computed within each segment to determine refined integrated ratings. After sorting the data according to new ratings, it’s further reduced to $\eta\%$ of its initial size. The number of segments is increased by a factor of $\beta$, unless it reaches the maximum threshold $n_{\text{max}}$, in which case the data is divided into $n_{\text{max}}$ segments. The iterative process of calculating orthogonality within progressively smaller subsets continues before the subsequent reduction leaves less than $k$ data points for selection.
@@ 4.2 Main Results
Comparing the FIRE results to those of the individual raters, our rating integration method demonstrates significant improvements, with an average score increase of up to 1.9% over the best single rater. Additionally, our method outperforms Random, QuRating and other data selection methods, validating the high quality of the data selected by FIRE.
@@
FIRE is superior to other integration methods.
When compared to other integration methods, FIRE proves to be more effective. As shown in Table 1, Comprehensive Rater with a multi-dimension prompt results in an average score that is even lower than the single-dimension rater Facts and Trivia. This suggests that GPT-4o still falls short in assessing data quality from a broad perspective. Both Mix Criteria and Max Criteria are inferior to FIRE, indicating that a comprehensive evaluation is more beneficial. Average simply calculates the mean of all ratings; however, different dimensions should not have an equal impact on model performance, and the experimental results of FIRE demonstrate an improvement over Average.
@@ 4.3 Analysis
Figure 4: The in-context learning results with respect to pretraining FLOPS on four downstream tasks: ARC-E, ARC-C, SciQ, and HellaSwag. FLOPS are approximated as 6 × model size × dataset size Kaplan et al. (2020)
Figures 1 and 4 show how the model’s performance on downstream tasks evolves with the pretraining FLOPS. In terms of average score, our method outperforms the random baseline by 2.9%. From the training efficiency perspective, our method reduces training FLOPS to achieve a certain performance level by more than half. In addition, our method shows a significant advantage in the ARC-E/C and SciQ tasks, consistently scoring higher than the random baseline. However, on the HellaSwag task, our method’s performance is similar to the random baseline and does not consistently surpass it. One possible explanation is that HellaSwag is an especially challenging dataset, which makes it difficult to discern performance differences on the 1.3B scale model.
@@
Method FLOPS ARC-C HellaSwag AVG.
@@
Figure 6: (a) The Pearson correlation between different raters. (b) Illustration of top samples rated by each method. We use Sentence-T5 to encode each text, and employ t-SNE to perform dimensionality reduction on dense vectors.
@@
Moreover, we analyze the data properties that FIRE focuses on compared to other raters. Based on the four dimensions in QuRating and a comprehensive evaluation dimension, we determine the percentage of high-quality data selected by each rater in each dimension. We pick 1000 data points at random from the top 20B data that are selected by each method. Then we use GPT-4o to assess each dimension six times, taking the average as the final evaluation result. Refer to the Appendix D.3 for more details. As we can see from Table 3: (1) From a comprehensive perspective, FIRE shows the best results, as evidenced by our performance in downstream tasks. (2) In terms of each dimension, FIRE consistently achieves relatively high accuracy, demonstrating that the data selected by this method maintains high quality across all dimensions.
Based on the results of previous experiments, compared to single raters, our method could potentially select more diverse samples, which is beneficial for model training. In order to validate this hypothesis, we take the top-rated 1000 texts from each method and encode these texts using Sentence-T5 Ni et al. (2022). Subsequently, we reduce dimensions using t-SNE Van der Maaten and Hinton (2008), with the visualization results displayed in Figure 6(b). In the latent semantic space, the top data selected by single raters tends to be clustered, whereas data selected by FIRE spans a broader range, ensuring the diversity of the selected data. Moreover, as seen from previous case studies, the data selected by FIRE also maintains high quality, suggesting that FIRE has achieved a good balance between diversity and quality.
@@ 5 Related Works
When pretraining language models, a large amount of text corpus is often crawled from the internet. However, several studies Li et al. (2023); Zhou et al. (2024) suggest that high-quality data is more beneficial to the model’s performance. To select high-quality data, a common strategy involves utilizing rules crafted by experts Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024) and removing duplicate sentences Lee et al. (2022); Sorscher et al. (2022); Abbas et al. (2023); Soboleva et al. (2023); Tirumala et al. (2024). However, they often fall short in effectively selecting high-quality data based on semantic content. An alternative approach involves utilizing a target data source or proxy model Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023); Thakkar et al. (2023); Engstrom et al. (2024); Yu et al. (2024).
Training a classifier is a more straightforward method Du et al. (2022); Gururangan et al. (2022); Zhang et al. (2024); Wettig et al. (2024); Sachdeva et al. (2024). Du et al. (2022) implemented a logistic regression binary classifier to score the data, while some studies train more complex scorers Zhang et al. (2024); Sachdeva et al. (2024). Additionally, QuRating Wettig et al. (2024) trains multiple raters with a finer-grained approach to analyze the contribution of data to model performance improvement from different dimensions.
The methods above primarily focus on selecting data from a single aspect. While QuRating has developed raters across multiple dimensions, it has not thoroughly explored how to integrate these raters, which is meticulously addressed in our research.
@@ 6 Conclusion
We propose FIRE, a flexible and scalable framework for integrating multiple data quality raters, enabling a comprehensive assessment of data quality across various dimensions. Initially, we align the ratings from different dimensions, mapping them to a unified score space. We then introduce orthogonality to adjust the weights of different raters. Given the variation in orthogonality across different rating ranks, we propose a progressive approach for fine-grained data selection. Experiments on the SlimPajama dataset reveal that FIRE outperforms other data selection methods, significantly enhancing the pretrained model’s performance across a wide range of downstream tasks. It achieves a 2.9% average performance improvement over Random and reduces the FLOPs required to reach a certain performance level by more than half.

### v3

@@ Abstract
Selecting high-quality data can improve the pretraining efficiency of large language models (LLMs). Existing methods generally rely on heuristic techniques or single quality signals, limiting their ability to evaluate data quality comprehensively. In this work, we propose FIRE, a flexible and scalable framework for integrating multiple data quality raters, which allows for a comprehensive assessment of data quality across various dimensions. FIRE aligns multiple quality signals into a unified space, and integrates diverse data quality raters to provide a comprehensive quality signal for each data point. Further, we introduce a progressive data selection scheme based on FIRE that iteratively refines the selection of high-quality data points. Extensive experiments show that FIRE outperforms other data selection methods and significantly boosts pretrained model performance across a wide range of downstream tasks, while requiring less than 37.5% of the training data needed by the Random baseline to reach the target performance.
@@
Liangyu Xu4,  Xuemiao Zhang1,4,  Feiyu Duan2,4thanks: Equal contribution.,  Sirui Wang3,4thanks: Corresponding author., Rongxiang Weng4,  Jingang Wang4,  Xunliang Cai4 1 Peking University 2 Beihang University 3 Tsinghua University 4 Meituan zhangxuemiao@pku.edu.cnduanfeiyu@buaa.edu.cn {xuliangyu02, wangsirui, wengrongxiang, wangjingang02, caixunliang}@meituan.com
@@ 1 Introduction
Figure 1: Downstream accuracy with respect to pretraining tokens for Random, FIRE, and FIRE Progressive.
@@
Extensive experiments demonstrate that by applying integrated ratings from multiple raters, our method achieves superior results across a variety of downstream tasks. Figure 1 illustrates that FIRE significantly enhances the pretrained model. We summarize our main contributions as follows:
@@
(3) Extensive experiments demonstrate that FIRE enhances the pretrained model’s performance by an average of 2.9%, while requiring less than 37.5% of the training data needed by the Random baseline to reach the target performance.
@@ 2.1 Overview of the Method
We propose FIRE, a method that flexibly integrates multiple raters to comprehensively evaluate data quality. It involves two key processes: (a) Rating Alignment and (b) Rater Integration. Figure 2 illustrates the overall framework of FIRE. Many off-the-shelf raters exist in practice. To be integrated by FIRE, a rater must provide a scalar score for each data point and be empirically validated for effectiveness.
@@
$I(x)=\mathbf{A}(x)^{T}(\mathbf{o}\odot\bm{\gamma})$ (1)
where $\mathbf{A}(x)$ is the vector of aligned ratings for data point $x$, $\mathbf{o}$ is the overall orthogonality vector, $\bm{\gamma}$ represents the intrinsic reliability vector of the raters, and $\odot$ denotes the element-wise product.
@@ 2.2 Rating Alignment
Uniformly sample a subset $\mathcal{D}_{r}$ from the pretraining data $\mathcal{D}_{t}$. It is worth noting that the data quality distributions of $\mathcal{D}{r}$ and $\mathcal{D}{t}$ are consistent. This consistency allows us to use $\mathcal{D}{r}$ as a representative reference set: if a subset of data is of higher quality than data in $\mathcal{D}{r}$, it can be considered to exceed the typical quality of $\mathcal{D}_{t}$, and vice versa.
@@
where $\mathcal{D}_{ij}$ signifies the subset sampled from the $j$-th rating interval for rater $R_{i}$. The win rate $w_{ij}$ is the proportion of samples $x$ in $\mathcal{D}_{ij}$ that GPT-4o determines have higher quality than the comparison sample $y$ from $\mathcal{D}_{r}$. Calculating win rates relative to $\mathcal{D}_{r}$ makes ratings from multiple raters comparable. Moreover, since we sample only 1,000 data points per interval for comparison, the computational cost of this process remains low. The prompt for GPT-4o is detailed in Appendix C.2. We demonstrate the reliability of using GPT-4o for quality comparison in Appendix C.3.
@@ 2.3 Rater Integration
Suppose for raters $R_{1},R_{2},\ldots,R_{n}$, each rater corresponds to a standard basis vector $\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}$ in the quality space. The integrated quality vector $\mathbf{q}(x)$ of data point $x$ can be expressed as:
$\mathbf{q}(x)=\sum_{j=1}^{n}\gamma_{j}A_{j}(x)\mathbf{v}_{j}$ (3)
where $\gamma_{j}$ denotes the intrinsic reliability of rater $j$, $A_{k}(x)$ denotes the rating for data point $x$ from rater $R_{k}$ after alignment. Ideally, if $\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}$ form an orthogonal basis, it is reasonable to measure the overall quality of the data using the L1 norm of $\mathbf{q}(x)$, as it represents the sum of the scores of data point $x$ across various orthogonal quality dimensions. However, $\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}$ are not necessarily completely independent. There may be raters $R_{i}$ and $R_{j}$ with a correlation coefficient $\rho>0$ and directly adding the corresponding aligned ratings would increase the weight of a particular quality dimension. To mitigate this issue, we define $O(i,j)$ to quantify the orthogonality of two raters $i$ and $j$ (the formalization of $O(i,j)$ can be found in Appendix C.6). For a rater $R_{i}$, we apply the sum of its orthogonality with all other raters to weight its rating. If a rater is highly correlated with others, we use a lower orthogonality to penalize. So the integrated rating for data point $x$ can expressed as
$I(x)=\sum_{j=1}^{n}\gamma_{j}o_{j}A_{j}(x)$ (4)
@@
where $d$ is the damping factor and $k$ denotes the $k$-th iteration. In PageRank, node centrality depends on edge weights. Since FIRE defines edge weights via orthogonality, the resulting scores reflect each node’s overall independence. Since the introduction of the damping factor aims to address the rank sinks problem, which is not present in our graph, we find it reasonable to set $d=1$. Assuming the number of iterations is $\alpha$, and normalizing the final result, our final formula becomes:
@@
Final version of the integrated rating for data point $x$ can be expressed as:
$I(x)=\mathbf{A}(x)^{T}(\mathbf{o}\odot\bm{\gamma})$ (9)
where $\mathbf{A}(x)=[A_{1}(x),A_{2}(x),\ldots,A_{n}(x)]^{T}$ is the vector of aligned ratings for data point $x$ from all raters, $\mathbf{o}$ is the overall orthogonality vector, $\bm{\gamma}=[\gamma_{1},\gamma_{2},\ldots,\gamma_{n}]^{T}$ represents the intrinsic reliability vector, and $\odot$ denotes the Hadamard product.
@@
12:     Select the top $\eta\%$ data according to $\mathcal{S}_{P_{i}}$
@@
3 Progressive Data Selection via FIRE
@@
Specifically, as shown in Algorithm 1, we first calculate the integrated ratings for the data points in $\mathcal{D}_{t}$ based on FIRE, then sort the data points, but we don’t select them right away based on these ratings. It is reduced to $\eta\%$ of its original size by selecting the top $\eta\%$ of data based on integrated ratings. Then, the data is partitioned into $n$ segments based on integrated ratings’ quantiles. Orthogonality is computed within each segment to determine refined integrated ratings. After sorting the data according to new ratings, it’s further reduced to $\eta\%$ of its initial size. The number of segments is increased by a factor of $\beta$, unless it reaches the maximum threshold $n_{\text{max}}$, in which case the data is divided into $n_{\text{max}}$ segments. The iterative process of calculating orthogonality within progressively smaller subsets continues before the subsequent reduction leaves less than $k$ data points for selection.
@@ 4.2 Main Results
FIRE is superior to other integration methods.
FIRE demonstrates greater effectiveness than existing integration methods, while introducing minimal additional computational cost (see Appendix D.3 for a detailed analysis of computational cost). As shown in Table 1, Comprehensive Rater with a multi-dimension prompt results in an average score that is even lower than the single-dimension rater Facts and Trivia. This suggests that GPT-4o still falls short in assessing data quality from a broad perspective. Both Mix Criteria and Max Criteria are inferior to FIRE, indicating that a comprehensive evaluation of FIRE is more beneficial. Average simply calculates the mean of all ratings and the experimental results of FIRE demonstrate an improvement over Average. To demonstrate the robustness of FIRE, we conduct additional experiments integrating other raters. Detailed results are provided in Appendix D.4.
@@
Comparing FIRE to the individual raters, it demonstrates significant improvements, with an average score increase of up to 1.9% over the best single rater and 2.9% over random selection. From a data efficiency perspective, FIRE achieves comparable performance using less than 37.5% of the data required by the Random baseline. Additionally, our method outperforms QuRating and other data selection methods, validating the high quality of the data selected by FIRE.
@@ 4.3 Analysis
Figure 4: The in-context learning results with respect to pretraining tokens on four downstream tasks: ARC-E, ARC-C, SciQ, and HellaSwag.
Figures 1 and 4 show how the model’s performance on downstream tasks evolves with the pretraining tokens. In terms of average score, our method outperforms the random baseline by 2.9%. From the training efficiency perspective, our method reduces training tokens to achieve a certain performance level by more than half. In addition, our method shows a significant advantage in the ARC-E/C and SciQ tasks, consistently scoring higher than the random baseline. However, on the HellaSwag task, our method’s performance is similar to the random baseline and does not consistently surpass it. One possible explanation is that HellaSwag is an especially challenging dataset, which makes it difficult to discern performance differences on the 1.3B scale model.
@@
Method Pretraining FLOPS ARC-C HellaSwag AVG.
@@
Figure 6: (a) The Pearson correlation between different raters. (b) Illustration of top samples rated by each method. We use Sentence-T5 to encode texts, and employ t-SNE to perform dimensionality reduction.
@@
Moreover, we analyze the data properties that FIRE focuses on compared to other raters. Based on the four dimensions in QuRating and a comprehensive dimension, we determine the percentage of high-quality data selected by each rater in each dimension. We pick 1000 data points at random from the top 20B data that are selected by each method. Then we use GPT-4o to assess each dimension six times, taking the average as the final evaluation result. Refer to the Appendix D.5 for more details. As shown in Table 3: (1) From a comprehensive perspective, FIRE shows the best results, as evidenced by our performance in downstream tasks. (2) In terms of each dimension, FIRE consistently achieves relatively high accuracy, demonstrating that the data selected by this method maintains high quality across all dimensions.
To assess whether FIRE selects more diverse samples than single raters, we extract the top 1000 texts from each method and encode them using Sentence-T5 Ni et al. (2022). We then apply t-SNE Van der Maaten and Hinton (2008) for dimensionality reduction and visualize the results in Figure 6(b). In the latent semantic space, samples selected by single raters appear more clustered, while those selected by FIRE are more broadly distributed, indicating higher diversity. Combined with prior case studies showing high quality, FIRE demonstrates a strong balance between diversity and quality.
@@ 5 Related Works
When pretraining language models, a large amount of text corpus is often crawled from the internet. However, several studies Li et al. (2023); Zhou et al. (2024); Duan et al. (2025) suggest that high-quality data is more beneficial to the model’s performance. To select high-quality data, a common strategy involves utilizing rules crafted by experts Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024) and removing duplicate sentences Lee et al. (2022); Sorscher et al. (2022); Abbas et al. (2023); Soboleva et al. (2023); Tirumala et al. (2024). However, they often fall short in effectively selecting high-quality data based on semantic content. An alternative approach involves utilizing a target data source or proxy model Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023); Thakkar et al. (2023); Engstrom et al. (2024); Yu et al. (2024).
Training a classifier is a more straightforward method Du et al. (2022); Gururangan et al. (2022); Zhang et al. (2024); Wettig et al. (2024); Sachdeva et al. (2024). Du et al. (2022) implemented a logistic regression binary classifier to score the data, while some studies train more complex scorers Zhang et al. (2024); Sachdeva et al. (2024). Additionally, QuRating Wettig et al. (2024) trains multiple raters with a finer-grained approach to analyze the contribution of data to model performance improvement from different dimensions. Other studiesZhang et al. (2025b, a) explore methods to boost pretraining efficiency by curriculum learning.
Prior methods mainly select data from a single perspective. Although QuRating introduces multidimensional raters, it does not systematically address their integration—a challenge our work explicitly tackles.
@@ 6 Conclusion
We propose FIRE, a flexible and scalable framework that integrates multiple data quality raters for comprehensive, multi-dimensional data assessment. First, ratings from different dimensions are aligned into a unified space. Then, orthogonality is introduced to adjust rater weights. To handle orthogonality variations across rating ranks, we adopt a progressive approach for fine-grained data selection. Experiments on the SlimPajama dataset show that FIRE outperforms other selection methods, substantially improving pretrained model performance across diverse downstream tasks.